OpenStudy (anonymous):

The volume in cubic feet of a box can be expressed as v(x)=x^3-6x^2+8x , or as the product of three linear factors with integer coefficients. The width of the box is 2 – x. Factor the polynomial to find linear expressions for the height and the width.

OpenStudy (ghazi):

x(x-4)(x-2)

OpenStudy (anonymous):

How do i factor that?

OpenStudy (ghazi):

factor of x^3-6x^2+8x

OpenStudy (ghazi):

x*(x^2-6x+8)....take x as common from the cubic equation

OpenStudy (anonymous):

How? I dont know what your talking about

OpenStudy (ghazi):

see in your given equation \[x^3+6x^2+8x= x*(x^2+6x+8)= x*(x-4)(x-2)\]

OpenStudy (eyust707):

Em, that \(\Rightarrow\) x(x-4)(x-2) is the factored form

OpenStudy (eyust707):

Now the equation for the volume of a cube is Voume=(Length)(Width)(Height)

OpenStudy (eyust707):

They tell us the x-2 part is the width

OpenStudy (anonymous):

Omg. I dont know.

OpenStudy (anonymous):

Nevermind.

OpenStudy (ghazi):

@EmilyJernigan done? :)

OpenStudy (anonymous):

No. I give up.

OpenStudy (ghazi):

he he :) wait

OpenStudy (eyust707):

Height = x Width = x-2 Length = x-4

OpenStudy (anonymous):

Thank you!